OpenStudy (anonymous):

Which of the following is the general form of the equation 5x^2 - 20x = 15?

OpenStudy (anonymous):

x^2 - 4x = 3 x^2 - 4x - 3 = 0 5x^2 - 20x - 15 = 0

OpenStudy (anonymous):

@ARMYRANGERS

OpenStudy (armyrangers):

Ok well witch one do think it is

OpenStudy (anonymous):

a

OpenStudy (armyrangers):

You forgot to set it equal to zero

OpenStudy (anonymous):

so its b? lol

OpenStudy (armyrangers):

Yes lol,